# Larger Is Not Always Better: Exploring Small Open-source Language Models in Logging Statement Generation

### v2

@@ 3.5.1. Tradictional evaluating metrics
Logging level: We use the Level Accuracy (LA) and Average Ordinal Distance Score (AOD) for evaluating the prediction of logging levels. Given the significant semantic differences between these levels and their implications for system monitoring and maintenance, we rigorously assess LA by comparing predicted log levels against their ground truth values in the source code. This metric is formally defined as the ratio of correctly predicted log levels ($LS_{level\_correct}$) to the total number of logging statements ($LS_{all}$), expressed as $\frac{LS_{level\_correct}}{LS_{all}}$. AOD evaluates how closely the current logging level aligns with the recommended logging level for each specific logging statement, as detailed in (Li et al., 2021a). The formula to calculate AOD is given by: $AOD=\frac{\sum_{i=1}^{N}(1-\frac{Dis(a_{i},s_{i})}{MaxDis(a_{i})})}{N}$, where $N$ represents the total number of logging statements in consideration. The term $MaxDis(a_{i})$ is used to denote the maximum potential distance for the actual log level $a_{i}$.
@@ 3.6. Implimentation Details
To evaluate conventional logging approaches and LLMs, we reproduced conventional methods using replication packages provided by their authors. For LLMs, we generated logging statements by calling their official APIs, setting the temperature to 0 to ensure deterministic outputs for identical queries, thus guaranteeing reproducibility. For SOLMs, we performed fine-tuning for one epoch on a dedicated dataset, using a batch size of 64 and a learning rate of 1e-4. All experiments, including training, fine-tuning, and inference, were conducted on a single NVIDIA A100 80GB GPU provided by Modal (Modal Labs, Inc., 2024), a serverless cloud infrastructure platform that supports easy deployment and reproducibility through provided scripts. Detailed experimental settings are available in our replication package (Renyi Zhong, Yichen Li, Guangba Yu, Wenwei Gu, Jinxi Kuang, Yintong Huo, Michael R. Lyu, 2025).
@@ 4.1. RQ1: What are the most effective prompting strategies for using SOLMs in automated logging?
Second, RAG emerges as the most effective prompting technique for enhancing automated logging generation genertaion performance. While other techniques occasionally yield the top score for an isolated metric, RAG demonstrates the most significant and robust improvements across the majority of metrics and models. Notably, for Mistral, RAG achieves the highest scores across all evaluated metrics, including PA (18.15), F1 (40.21), and ROUGE-L (34.74). For LLaMA3, RAG secures the top performance in PA (21.01), PMR (41.68), F1 (46.28), BLEU-4 (15.95), and ROUGE-L (36.73). For CodeLlama and Qwen2.5-coder, RAG generally leads to substantial gains over the baseline, particularly in metrics like F1 (CodeLlama: 39.01, Qwen2.5-coder: 49.28) and ROUGE-L (CodeLlama: 34.36, Qwen2.5-coder: 37.57). These results strongly indicate that providing relevant contextual information retrieved from a knowledge base significantly aids the SOLMs in accurately determining where to log, what variables to include, and formulating appropriate log messages, making RAG a highly promising strategy.
@@
Table 2. Comparison of prompting techniques for automated logging generation using four 7B instruction-following SOLMs.
@@ 4.2. RQ2: What’s the best tuning strategy using SOLMs for automated logging?
(RQ2.3) Does the instruct variant of a SOLM outperform its base counterpart for automated logging? The distinction between using a ‘base’ pre-trained model versus an ‘instruct’ version of an SOLM presents another critical strategic choice. Instruct models are tuned to follow user prompts and formats, which is useful for specific tasks, but their tuning is generally broad. Base models, on the other hand, represent the raw capabilities learned during pre-training and might offer greater plasticity when fine-tuned on a highly specific downstream task like ours. Therefore, we aim to clarify which model mode serves as a better foundation for this task.
@@
To address RQ2.3, we compare the performance of ‘base’ models against their ‘instruct’ counterparts for automated logging. We select the Mistral 7B model, available in both base and instruct variants, for a controlled comparison, as it got the best perfomance in RQ2.1. Both model modes are fine-tuned using LoRA and employed the RAG-enhanced prompt, consistent with the result of RQ2.1 and RQ2.2. The fine-tuning process spanned five epochs, and performance are evaluated using metrics including Reject Rate, PA, LA, AOD, PMR, F1, BLEU-4, and ROUGE-L. The Reject Rate, defined as the percentage of test set examples rejected by the fine-tuned model during inference due to misalignment with learned task-specific criteria, quantifies the model’s selectivity, reflecting its ability to adhere to prompt instructions and generate valid logging statements.
Table 3. Performance Comparison of PEFT Techniques for four 7B SOLMs in automated logging.
@@
Additionally, the venn diagram in Figure 4(a) illustrates the overlap of corrected logging statement placements across different PEFT variations for Mistral-7B. The largest overlap (183) is observed in the central region, indicating a core set of logging statements consistently corrected across all PEFT methods (base, Prefix Tuning, Prompt Tuning, LoRA, and QLoRA). LoRA and QLoRA show significant individual contributions (125 and 86, respectively), suggesting their effectiveness in identifying unique logging placements, while the base method contributes the least (40), highlighting the improvement brought by PEFT techniques.
@@
Figure 4. Overlap of Corrected Logging Statement Placement Across Different SOLM Configurations.
Table 4. Impact of Model Size on Performance and Resource Usage for automated logging.
@@
The venn diagram in Figure 4(b) depicts the overlap of corrected logging statement placements across different sizes of the Qwen2.5Coder model (0.5B, 1.5B, 3B, 7B, 14B). The central overlap (442) represents logging statements consistently corrected across all sizes, with the 14B model contributing the most unique placements (152), followed by 7B (73). Smaller models (0.5B and 1.5B) show limited unique contributions (21 and 6, respectively), reinforcing the finding that models with 3B+ parameters perform better for this task.
@@
The venn diagram in Figure 4(c) further illustrates the overlap of corrected logging statement placements across epochs for the Mistral-7B-Instruct model. The central overlap (1120) indicates a stable core of corrections maintained across all five epochs, with the first epoch contributing the most unique corrected placements (40). The decline in unique contributions from later epochs (e.g., 3 for the fifth epoch) supports the observation of diminishing returns and potential overfitting with prolonged fine-tuning.
@@ 4.3. RQ3: How effectively do SOLMs compare to existing methods and LLM baselines in automated logging?
(RQ3.2) For statement quality, Qwen2.5-coder-14B-RAG-LoRA again leads across all metrics. These results surpass the best-performing LLM, deepseek-coder-v3-RAG (e.g., 68.39% LA, 18.97% BLEU-4), and existing methods like Fastlog (e.g., 59.26% LA, 13.28% BLEU-4). CodeLlama-13B-RAG-LoRA also performs strongly, particularly in BLEU-4 (24.39%) and ROUGE-L (46.91%), closely rivaling Qwen2.5-coder-14B. The superior performance of fine-tuned SOLMs suggests that targeted optimization enables them to generate more accurate, relevant, and contextually appropriate logging statements than both unoptimized LLMs and traditional methods.
Figure 7 presents a case study comparing logging statements generated by different models for an error logging scenario in a Java method handling a DuplicateKeyException. The ground truth logging statement is log.error(”Update alert group error, groupName:{}”), alertGroup.getGroupName(), ex). Among the LLMs, Claude3.7sonnet-RAG generates a statement that closely resembles the ground truth but includes an additional variable (id), which may add unnecessary verbosity. Deepseek-coder-v3-RAG and GPT4o-RAG produce statements with missing or misused variables (e.g., using groupName directly instead of alertGroup.getGroupName()), reducing their contextual accuracy. LLaMA3.1-405B-RAG omits the exception variable (ex), limiting its diagnostic utility. In contrast, fine-tuned SOLMs demonstrate superior performance: Mistral-7B-RAG-LoRA and CodeLlama-13B-RAG-LoRA generate statements identical to the ground truth, ensuring both accuracy and relevance. Qwen2.5-coder-14B-RAG-LoRA adds minor additional text (”already exist”) but retains all critical components, closely aligning with the ground truth.
Table 5. Performance Comparison of Automated Logging Approaches Across All Metrics.
@@
Large Language Model
@@
Fine-tuned Small Open-source Language Models
@@
Figure 7. A Case of Generated Logging Statement from Multiple Models.
@@
Findings 7: Fine-tuned SOLMs outperform both existing methods and LLMs across all evaluated metrics, demonstrating superior logging location accuracy and statement quality.
4.4. RQ4: Can SOLMs generalize logging statement generation across diverse code repositories?
Motivation. In real-world software development, logging practices vary significantly across projects due to differences in coding styles, project domains, and developer preferences. For SOLMs to be practically viable for automated logging, they must demonstrate robust generalization when applied to unseen repositories. Poor generalization could lead to ineffective logging statements that fail to capture critical runtime information or adhere to project-specific conventions, limiting the utility of SOLMs in cross-project settings.
@@ 4.4. RQ4: Can SOLMs generalize logging statement generation across diverse code repositories?
Approach. To evaluate the cross-repository generalization ability of SOLMs, we design an experiment that trains models on a subset of repositories from the dataset and tests them on a distinct, non-overlapping set of repositories. We randomly partition the dataset into two groups, each containing five repositories, ensuring diversity in project domains. In each experimental run, we use one group of five repositories as the training set, three repositories from the other group as the validation set, and the remaining two repositories from the same group as the test set. The validation set primarily served to support RAG by providing a pool of examples from which the most similar code snippets are retrieved using the BM25 algorithm. This partitioning ensures that the test set represents unseen projects with potentially different coding styles and logging conventions, simulating real-world cross-repository scenarios. The experimental setup involves training on a subset of repositories (R1, R3, R5, R7, R9). It also includes validating on another subset (R2, R4, R6) and testing on unseen repositories (R8, R10) in the first configuration. The second configuration, meanwhile, trains on Apache-dominated repositories (R1, R3, R5, R7, R9) and tests on non-Apache repositories. For this experiment, we select the two best-performing 7B SOLMs from RQ2.1: Mistral and Qwen2.5-coder. We fine-tune these models using the best-performing strategy identified in prior experiments, combining LoRA with RAG.
@@
Results. Table 6 presents the performance of two fine-tuned SOLMs, Mistral-7B and Qwen2.5-coder-7B, in automated logging across diverse code repositories. The results demonstrate the generalization capabilities of SOLMs and highlight the impact of similar logging practices on cross-project performance.
@@
Findings 8: SOLMs demonstrate strong generalization capabilities in automated logging, maintaining high performance across diverse, unseen repositories.
@@
Findings 9: Fine-tuning SOLMs with data reflecting similar logging practices, such as those prevalent in Apache open-source projects, significantly enhances their cross-project generalization capabilities.

### v3

@@ 3.5.1. Tradictional evaluating metrics
Logging level: We use the Level Accuracy (LA) and Average Ordinal Distance Score (AOD) for evaluating the prediction of logging levels. Given the significant semantic differences between these levels and their implications for system monitoring and maintenance, we rigorously assess LA by comparing predicted log levels against their ground truth values in the source code. This metric is formally defined as the ratio of correctly predicted log levels ($LS_{level\_correct}$) to the total number of logging statements ($LS_{all}$), expressed as $\frac{LS_{level\_correct}}{LS_{all}}$. AOD evaluates how closely the current logging level aligns with the recommended logging level for each specific logging statement, as detailed in (Li et al., 2021a). The formula to calculate AOD is given by: $AOD=\frac{\sum_{i=1}^{N}(1-\frac{Dis(a_{i},s_{i})}{MaxDis(a_{i})})}{N}$, where $N$ represents the total number of logging statements in consideration. The term $MaxDis(a_{i})$ is used to denote the maximum potential distance for the actual log level $a_{i}$. We verify that all 10 projects within our dataset consistently utilize the same five logging levels (trace, debug, info, warn, and error), making a unified distance appropriate for this AOD metric.
@@ 3.6. Implimentation Details
To evaluate conventional logging approaches and LLMs, we reproduced conventional methods using replication packages provided by their authors. For LLMs, we generated logging statements by calling their official APIs. To promote deterministic outputs, we set the temperature to 0. We acknowledge that for some proprietary services like OpenAI, setting the temperature to 0 alone does not guarantee reproducibility (OpenAI, 2025).
For the fine-tuning of SOLMs, we employed the Axolotl framework. The training objective was a standard causal language modeling task, where the model learns to predict the next token in a sequence. We utilized Cross-Entropy Loss as the objective function to minimize during training. The training data was carefully prepared to align with our prompting strategies. For each data instance, the input source code was formatted using a specific prompt template (e.g., Base, RAG, or CoT) to create the final input sequence. The ground-truth code containing the logging statement served as the target label. All SOLMs were fine-tuned for one epoch. We use a learning rate of 1e-4 with a cosine scheduler, a global batch size of 64, and the Adam optimizer. For LoRA and QLoRA, we used a rank (r) of 16 and an alpha of 32. All experiments, including training, fine-tuning, and inference, were conducted on a single NVIDIA A100 80GB GPU provided by Modal (Modal Labs, Inc., 2024), a serverless cloud infrastructure platform. To ensure full reproducibility, the complete scripts, configuration files, and data preparation details are publicly available in our replication package (Renyi Zhong, Yichen Li, Guangba Yu, Wenwei Gu, Jinxi Kuang, Yintong Huo, Michael R. Lyu, 2025).
@@ 4.1. RQ1: What are the most effective prompting strategies for using SOLMs in automated logging?
Second, RAG emerges as the most effective prompting technique for enhancing automated logging generation performance. While other techniques occasionally yield the top score for an isolated metric, RAG demonstrates the most significant and robust improvements across the majority of metrics and models. Notably, for Mistral, RAG achieves the highest scores across all evaluated metrics, including PA (18.15), F1 (40.21), and ROUGE-L (34.74). For LLaMA3, RAG secures the top performance in PA (21.01), PMR (41.68), F1 (46.28), BLEU-4 (15.95), and ROUGE-L (36.73). For CodeLlama and Qwen2.5-coder, RAG generally leads to substantial gains over the baseline, particularly in metrics like F1 (CodeLlama: 39.01, Qwen2.5-coder: 49.28) and ROUGE-L (CodeLlama: 34.36, Qwen2.5-coder: 37.57). These results strongly indicate that providing relevant contextual information retrieved from a knowledge base significantly aids the SOLMs in accurately determining where to log, what variables to include, and formulating appropriate log messages, making RAG a highly promising strategy.
@@
Table 3. Comparison of prompting techniques for automated logging generation using four 7B instruction-following SOLMs. The greener, the better.
@@ 4.2. RQ2: What’s the best tuning strategy using SOLMs for automated logging?
(RQ2.3) Does the instruct variant of a SOLM outperform its base counterpart for automated logging? The distinction between using a ‘base’ pre-trained model versus an ‘instruct’ version of an SOLM presents another critical strategic choice. The ‘instruct’ variants of SOLMs have been pre-tuned by their creators on a wide array of tasks to enhance their ability to follow general user prompts. In contrast, ‘base’ models are not tuned for instruction-following and represent the direct output of the pre-training phase. This presents a critical choice: Is it more effective to start with a model already skilled in following general instructions, or does the base model offer greater plasticity for specialization when fine-tuned on a narrow, specific task such as automated logging? Therefore, we aim to clarify which model mode serves as a better foundation for our task.
(RQ2.4) Which prompting strategy is most effective when combined with fine-tuned SOLMs? Our prior findings established LoRA as the optimal fine-tuning method (RQ2.1) and identified RAG as the most effective prompting strategy for base models (RQ1). However, since fine-tuning fundamentally alters a model’s behavior, we cannot assume that the best prompting strategy for a base model remains optimal for its fine-tuned counterpart. To build a methodologically sound basis for our final model evaluation in RQ3, we empirically determine which prompting strategy works most effectively in synergy with a LoRA-tuned model. This step is crucial to validate our choice of the definitive best-performing configuration.
@@
To address RQ2.3, we compare the performance of ‘base’ models against their ‘instruct’ counterparts for automated logging. We select the Mistral 7B model, available in both base and instruct variants, for a controlled comparison, as it got the best perfomance in RQ2.1. Both model modes are fine-tuned using LoRA and employed the RAG-enhanced prompt, consistent with the result of RQ2.1 and RQ2.2. The fine-tuning process spanned five epochs, and performance are evaluated using metrics including PA, LA, AOD, PMR, F1, BLEU-4, and ROUGE-L. To specifically assess the models’ adherence to instructions in this comparison, we also introduce a diagnostic metric: the Reject Rate. The Reject Rate is defined as the percentage of test set examples rejected by the fine-tuned model during inference due to misalignment with learned task-specific criteria; this quantifies the model’s selectivity, reflecting its ability to adhere to prompt instructions and generate valid logging statements.
To address RQ2.4, we conduct a controlled experiment applying all four prompting strategies to the LoRA-tuned versions of Qwen2.5-coder-7B. We evaluated the performance of four configurations (e.g., LoRA+ICL) on the test set. This comparative analysis allows us to isolate the impact of each prompting technique on an already specialized model and identify the most effective combination of fine-tuning and prompting for automated logging.
Table 4. Performance Comparison of PEFT Techniques for four 7B SOLMs in automated logging. The greener the better.
@@
Additionally, the venn diagram in Figure 4(a) illustrates the overlap of correctly predicted logging statement placements across different PEFT variations for Mistral-7B. The largest overlap (183) is observed in the central region, indicating a core set of logging statements whose placement were successfully identified by all PEFT methods . LoRA and QLoRA show significant individual contributions (125 and 86, respectively), suggesting their effectiveness in identifying unique logging placements, while the base method contributes the least (40), highlighting the improvement brought by PEFT techniques.
@@
Figure 4. Overlap of Correctly Logging Statement Placement Across Different SOLM Configurations.
Table 5. Impact of Model Size on Performance and Resource Usage for automated logging. The greener, the better.
@@
The venn diagram in Figure 4(b) depicts the overlap of correctly predicted logging statement placements across different sizes of the Qwen2.5-coder model (0.5B, 1.5B, 3B, 7B, 14B). The central overlap (442) represents logging statements whose placements were successfully identified across all model sizes, with the 14B model contributing the most unique placements (152), followed by 7B (73). Smaller models (0.5B and 1.5B) show limited unique contributions (21 and 6, respectively), reinforcing the finding that models with 3B+ parameters perform better for this task.
@@
The venn diagram in Figure 4(c) further illustrates the overlap of correctly predicted logging statement placements across epochs for the Mistral-7B-Instruct model. The central overlap (1120) indicates a stable core of accurately placed statements that were consistently identified across all five epochs, with the first epoch contributing the most unique correctly predicted placements (40). The decline in unique contributions from later epochs (e.g., 3 for the fifth epoch) supports the observation of diminishing returns and potential overfitting with prolonged fine-tuning.
@@
Table 6. Performance Comparison of Different Prompting Strategies Combined with LoRA.
Strategy Location Level Variable Text
PA LA AOD PMR F1 BLEU-4 ROUGE-L
LoRA+base 57.81 68.30 88.01 51.96 58.33 20.63 42.69
LoRA+ICL 60.77 66.08 87.59 51.45 57.75 21.39 44.06
LoRA+RAG 62.40 68.46 88.82 52.24 58.98 23.04 44.96
LoRA+COT 58.81 68.09 88.73 50.91 57.39 21.30 43.36
(RQ2.4) Table 6 presents the performance comparison of different prompting strategies when applied to the LoRA-tuned Qwen2.5-coder-7B model. The results clearly indicate that the combination of LoRA and RAG yields the most effective performance across all evaluated metrics. This configuration achieved the highest scores in position accuracy (i.e., PA: 62.40%), level accuracy (i.e., LA: 68.46%), variable prediction (i.e., F1: 58.98%), and text generation quality (i.e., BLEU-4: 23.04, ROUGE-L: 44.96%).
While other prompting strategies also benefited from the LoRA fine-tuning, their performance was consistently surpassed by the RAG-enhanced approach. For instance, LoRA+ICL, the second-best strategy, achieved a PA of 60.77%, which is 1.63 percentage points lower than LoRA+RAG. This outcome validates the hypothesis that the benefits of LoRA and RAG are synergistic. Fine-tuning with LoRA adapts the model to the specific structural and syntactic nuances of the automated logging task, while RAG provides critical, instance-specific context at inference time. This combination allows the specialized model to leverage the most relevant examples, leading to a significant performance boost. Therefore, we confirm that LoRA fine-tuning combined with RAG prompting is the optimal strategy for this task.
Findings 7: The RAG prompting strategy is most effective when combined with a LoRA-tuned model, outperforming other prompting techniques across all metrics. This confirms a synergistic effect between task-specific fine-tuning and context-aware prompting.
@@ 4.3. RQ3: How effectively do SOLMs compare to existing methods and LLM baselines in automated logging?
Table 7. Performance Comparison of Automated Logging Approaches Across All Metrics. The greener, the better.
@@
SCLogger 44.39 68.02 88.56 50.54 59.17 22.48 44.60
Proprietary Large Language Model
@@
Fine-tuned Small Open-source Language Model
@@
(RQ3.2) For statement quality, fine-tuned SOLMs demonstrate exceptional performance, outperforming both proprietary LLMs and all existing methods, including the strong SCLogger baseline. Qwen2.5-coder-14B-RAG-LORA leads on several key metrics, including the highest LA (69.92%), AOD (89.36%), and F1 score (59.93%). Notably, these results surpass not only the best-performing LLM, deepseek-coder-v3-RAG (e.g., 68.39% LA, 18.97% BLEU-4), but also the highly competitive SCLogger (e.g., 68.02% LA, 59.17% F1).
Furthermore, CodeLlama-13B-RAG-LORA excels in text generation quality, achieving the highest BLEU-4 (24.39) and ROUGE-L (46.91) scores. This performance is superior to that of SCLogger, which itself sets a high bar with a BLEU-4 of 22.48 and a ROUGE-L of 44.60. The ability of fine-tuned SOLMs to outperform such a strong, specialized tool underscores the effectiveness of SOLMs. This suggests that targeted optimization enables SOLMs to generate more accurate, relevant, and contextually appropriate logging statements than both proprietary LLMs and state-of-the-art specialized tools.
Table 8. Correct Predictions Considering the Three-dimensional Challenges of Automated Logging.
Position
Level
Variable
Proprietary Large Language Model
Claude3.7sonnet-RAG 34.10 13.22 22.51 9.26 20.91
GPT4o-RAG 44.22 12.22 17.75 8.26 17.55
Deepseek-coder-v3-RAG 34.80 12.99 20.08 7.63 24.51
LLaMA3.1-405B-RAG 44.96 10.99 15.95 9.02 19.08
Fine-tuned Small Open-source Language Models
LLaMA3-8B-RAG-LoRA 43.16 11.69 16.65 9.02 19.48
Mistral-7B-RAG-LoRA 36.03 10.92 19.28 8.59 25.17
CodeLlama-13B-RAG-LoRA 36.43 11.72 17.58 10.89 25.38
Qwen2.5-coder-14B-RAG-LoRA 33.80 11.56 18.55 8.36 27.74
To further address the relationship between metrics, we analyze the proportion of logging statements that are simultaneously correctly generated across the key dimensions of location, level, and variables. Table 8 presents this multi-dimensional evaluation. The results reveal that a significant portion of generated logs fail on at least one dimension, underscoring the task’s complexity. However, the fine-tuned SOLMs consistently produce a higher percentage of fully correct statements (correct position, level, and variables). Qwen2.5-coder-14B-RAG-LoRA again demonstrates its superiority, generating fully correct logs in 27.74% of cases, the highest of any model. This is closely followed by CodeLlama-13B (25.38%) and Mistral-7B (25.17%). Notably, these SOLMs outperform the best proprietary LLM, Deepseek-coder-v3-RAG (24.51%). This analysis reinforces that SOLMs excel not only on individual metrics but also in producing more integrally correct and practically useful logging statements.
@@
Findings 8: Fine-tuned SOLMs outperform both existing methods and LLMs across all evaluated metrics, demonstrating superior logging location accuracy and statement quality.
4.4. RQ4: How robust are SOLMs in handling diverse real-world scenarios for logging statement generation?
Motivation. In real-world software development, automated logging tools must demonstrate robust performance across diverse challenging scenarios beyond controlled experimental settings. Two critical dimensions of robustness are essential for practical deployment:
(RQ4.1) Can SOLMs generalize logging statement generation across diverse code repositories? Software projects exhibit significant heterogeneity in logging practices due to team preferences, domain requirements, and coding conventions. Understanding whether SOLMs can maintain effectiveness when deployed on unseen codebases with different logging practices is crucial for determining their practical viability as general-purpose logging tools.
(RQ4.2) How does code length affect SOLM performance in automated logging? Production codebases contain methods ranging from simple functions to complex implementations spanning hundreds of lines. Evaluating performance scaling with code length is essential to validate whether our findings about SOLM effectiveness hold across the full spectrum of code complexity.
@@ 4.4. RQ4: How robust are SOLMs in handling diverse real-world scenarios for logging statement generation?
Approach. (RQ4.1) To evaluate the cross-repository generalization ability of SOLMs, we design an experiment that trains models on a subset of repositories from the dataset and tests them on a distinct, non-overlapping set of repositories. We randomly partition the dataset into two groups, each containing five repositories, ensuring diversity in project domains. In each experimental run, we use one group of five repositories as the training set, three repositories from the other group as the validation set, and the remaining two repositories from the same group as the test set. The validation set primarily served to support RAG by providing a pool of examples from which the most similar code snippets are retrieved using the BM25 algorithm. This partitioning ensures that the test set represents unseen projects with potentially different coding styles and logging conventions, simulating real-world cross-repository scenarios.
The experimental setup involves training on a subset of repositories (R1, R3, R5, R7, R9). It also includes validating on another subset (R2, R4, R6) and testing on unseen repositories (R8, R10) in the first configuration. The second configuration, meanwhile, trains on Apache-dominated repositories (R1, R3, R5, R7, R9) and tests on non-Apache repositories. For this experiment, we select the two best-performing 7B SOLMs from RQ2.1: Mistral and Qwen2.5-coder. We fine-tune these models using the best-performing strategy identified in prior experiments, combining LoRA with RAG.
(RQ4.2) To comprehensively understand how code length impacts automated logging performance, we construct two test sets from the AL-Bench dataset: Test-Short containing code snippets with less than 512 tokens and Test-Long containing snippets exceeding 512 tokens. This division enables us to analyze whether model performance and the effectiveness of different strategies vary with code complexity, as longer code segments typically contain more intricate control flows and multiple logging-worthy events.
We evaluate all key configurations explored in previous RQs on both test sets, including the four prompting strategies (base, ICL, RAG, CoT) with Qwen2.5Coder-7B models with all PEFT techniques (Prefix Tuning, Prompt Tuning, LoRA, QLoRA). We also include a representative LLM baselines (i.e., Deepseek-coder-v3) for comprehensive comparison.
@@
Results. (RQ4.1) Table 9 presents the performance of two fine-tuned SOLMs, Mistral-7B and Qwen2.5-coder-7B, in automated logging across diverse code repositories. The results demonstrate the generalization capabilities of SOLMs and highlight the impact of similar logging practices on cross-project performance.
@@
Findings 9: SOLMs demonstrate strong generalization capabilities in automated logging, maintaining high performance across diverse, unseen repositories.
@@
Findings 10: Fine-tuning SOLMs with data reflecting similar logging practices, such as those prevalent in Apache open-source projects, significantly enhances their cross-project generalization capabilities.
Table 10. Performance Comparison of Different Prompt Strategies and PEFT Techniques on Different Length Code Datasets.
Strategy Dataset PA LA PMR ROUGE-L
Different Prompting Techniques
base Short 4.30 58.91 27.91 27.13
Long 0.80 87.50 16.67 29.44
$\Delta$ $\downarrow$3.50 $\uparrow$28.59 $\downarrow$11.24 $\uparrow$2.31
base+ICL Short 2.03 45.90 36.07 31.21
Long 0.63 63.16 21.05 28.56
$\Delta$ $\downarrow$1.40 $\uparrow$17.26 $\downarrow$15.02 $\downarrow$2.65
base+RAG Short 3.40 59.80 45.10 15.10
Long 0.27 100.00 50.00 43.75
$\Delta$ $\downarrow$3.13 $\uparrow$40.20 $\uparrow$4.90 $\uparrow$28.65
base+COT Short 13.25 51.01 20.35 10.58
Long 6.30 77.78 16.93 32.96
$\Delta$ $\downarrow$6.95 $\uparrow$26.77 $\downarrow$3.42 $\uparrow$22.38
Different PEFT Techniques
prefix tuning+RAG Short 25.27 65.48 44.14 35.47
Long 11.25 72.58 45.92 39.36
$\Delta$ $\downarrow$14.02 $\uparrow$7.10 $\uparrow$1.78 $\uparrow$3.89
prompt tuning+RAG Short 30.04 67.63 46.12 38.74
Long 13.86 74.24 48.55 42.30
$\Delta$ $\downarrow$16.18 $\uparrow$6.61 $\uparrow$2.43 $\uparrow$3.56
LoRA+RAG Short 62.40 68.46 52.24 44.96
Long 50.37 82.70 56.97 55.86
$\Delta$ $\downarrow$12.03 $\uparrow$14.24 $\uparrow$4.73 $\uparrow$10.90
QLora+RAG Short 60.21 67.87 51.55 44.03
Long 46.17 82.55 57.71 56.02
$\Delta$ $\downarrow$14.04 $\uparrow$14.68 $\uparrow$6.16 $\uparrow$11.99
(RQ4.2) Table 10 presents a comprehensive analysis of how code length affects SOLM performance across different prompting and PEFT strategies using Qwen2.5Coder-7B. For prompting techniques without fine-tuning, position accuracy consistently decreases on longer code while logging level accuracy improves substantially. The base strategy shows PA dropping from 4.30% to 0.80%, yet LA increases from 58.91% to 87.50%. RAG exhibits the most extreme pattern with PA plummeting to 0.27% but achieving perfect LA of 100% on long code. COT demonstrates the most balanced degradation with PA declining from 13.25% to 6.30% while LA improves from 51.01% to 77.78%. This suggests that longer code provides richer semantic context for level determination but severely complicates precise statement positioning.
PEFT techniques show superior resilience to code length increases. LoRA combined with RAG maintains the strongest position accuracy on long code at 50.37% while achieving substantial LA improvement from 68.46% to 82.70%. QLoRA demonstrates comparable robustness with 46.17% PA on long code and LA increasing from 67.87% to 82.55%. The lighter fine-tuning approaches, prefix and prompt tuning, show moderate resilience with PA values of 11.25% and 13.86% respectively on long code, significantly outperforming prompting-only approaches.
Counterintuitively, variable matching performance and text quality consistently improve on longer code across all configurations. Most strategies achieve better PMR on long code, with LoRA showing improvement from 52.24% to 56.97%, indicating that extended context helps models make more informed variable selection decisions. Similarly, ROUGE-L scores universally increase, with LoRA improving from 44.96% to 55.86%, demonstrating that additional contextual information enhances text generation quality despite positioning challenges.
Findings 11: Code length creates contrasting effects: position accuracy decreases while logging level accuracy, variable matching, and text quality improve, indicating that longer code provides richer semantic context but increases positioning complexity.
Across all configurations, PEFT techniques demonstrate superior resilience compared to prompting-only approaches. Notably, the performance hierarchy established in our earlier experiments remains consistent on longer code: LoRA+RAG maintains the highest position accuracy at 50.37%. This consistent ranking across different code lengths reinforces the effectiveness of our identified optimal strategies, with LoRA’s superior parameter efficiency proving particularly valuable when processing complex, lengthy code segments.
Findings 12: The relative performance hierarchy of different strategies remains consistent across code lengths, with LoRA+RAG maintaining its superiority over other PEFT techniques.
Table 11. Performance Comparison of Representative SOLM and LLM on Different Length Code Datasets.
Model Dataset PA LA PMR ROUGE-L
Deepseek-coder-v3-RAG Short 65.20 68.39 49.28 41.31
Long 55.23 75.98 50.57 46.88
$\Delta$ $\downarrow$9.97 $\uparrow$7.59 $\uparrow$1.29 $\uparrow$5.57
Qwen2.5-coder-14B-RAG-LoRA Short 66.20 69.92 54.53 46.51
Long 52.77 83.01 60.73 58.69
$\Delta$ $\downarrow$13.43 $\uparrow$13.09 $\uparrow$6.20 $\uparrow$12.18
Table 11 shows a direct comparison between the best fine-tuned SOLM and the leading LLM baseline. Notably, the fine-tuned SOLM, Qwen2.5-coder, experiences a more pronounced drop in position accuracy ($\downarrow$13.43) than the LLM baseline ($\downarrow$9.97). This suggests that the SOLM’s smaller model scale may face greater challenges in pinpointing the exact insertion location within the expanded and more complex context of long code.
Conversely, this comparison highlights the distinct effectiveness of fine-tuning for content generation. The SOLM outperforms the LLM in its ability to capitalize on the richer context to improve statement quality. For instance, Qwen2.5-coder achieves more improvement in its ROUGE-L score. ($\uparrow$12.18 vs. $\uparrow$5.57) and demonstrates similarly superior gains in both LA and PMR. This indicates that while model scale might influence long-range positional awareness, task-specific fine-tuning is highly effective at teaching the model what to log, enabling it to generate more accurate and contextually relevant content when more information is available.
Findings 13: While fine-tuned SOLMs show a sharper decline in position accuracy on long code than LLMs, potentially due to their smaller model scale, they achieve substantially greater improvements in content generation quality. This highlights that fine-tuning effectively trains SOLMs to leverage rich context for generating logging statements.